OpenStudy (sphott51):

Simplify (2 +3i) / (2-i) the answer is (1 +8i) / (5) Needs to know how to get there

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

\[(a+bi)(a-bi)=a^2+b^2\\ (2-i)(2-i)=2^2+1^2=5\]

OpenStudy (misty1212):

oops i meant \[(2-i)(2+i)=2^2+1^2=5\]

OpenStudy (misty1212):

so multiply \[\frac{2+3i}{2-i}\times \frac{2+i}{2+i}=\frac{(2+3i)(2+i)}{5}\]

OpenStudy (sphott51):

I foil the numerator?

OpenStudy (anonymous):

Yes! Foil the numerator, then simplify. Remember \[i ^{2}=-1\]

OpenStudy (sphott51):

Thanks